# LoX: Low-Rank Extrapolation Robustifies LLM Safety Against Fine-tuning

### v2

@@ Appendix C Effecive Rank Experiments
In this section, we show the experiments for determining the effective rank used in the experiments. To determine the effective rank that is essential for preserving safety (0% ASR), we perform ASR evaluation on the aligned model, by removing all ranks except the top-$k$ ones. The evaluated weights are given by the following expression.
@@
Figure 7: Evaluating ASR by preserving top-$k$ ranks. For most models, rank 3 is effective for achieving 0% ASR.
@@
As observed in Fig. 7, the safety can be recovered by preserving just a few ranks, in all cases. For the Llama model aligned with 65.6k examples, the effective rank is 6, while the other models the effective rank is 3.
@@ Appendix D Extended Robustness Metric Results
In Table 3, we present individual value results for Fig. 2 . It is possible to osberve that (a) robustness increases as the data size increases and (b) $R_{\text{ft}}<R_{\text{align}}$, indicating that the proportion of safety knowledge decreases after fine-tuning.
@@ Appendix E Extended Identity Shifting Attack Results
Table 4, Table 5, and Table 6 present individual run results for the experiments in Table 1, Table 2, and Fig. 5, respectively.
Table 4: ASR of individual Identity Shifting attack runs in correspondence to Table 1.
@@
Table 5: ASR of individual Identity Shifting attack runs in correspondence to Table 2.
@@
Table 6: ASR of individual Identity Shifting attack runs in correspondence to Fig. 5.
@@ Appendix F Broken Output Examples
To further showcase the model degradation effects of ExPO when compared to LoX, and quantitatively justify the observation in Section 5.2, we provide evaluation of the highest $\alpha$ models plotted on Fig. 5, and extrapolating further (which we considered broken), following (Lin et al., 2023). Table 7 shows that extrapolating full rank to $\alpha=1.25$ causes the largest degradation according to all metrics. In Fig. 8, we also present examples of broken and normal completions generated by these models.
Table 7: Utility metrics for $k=6$, full rank, and varying $\alpha$. For both values of $k$, we show the highest $\alpha$ from Fig. 5 and extrapolating further (which we consider broken). The lowest values for each $k$ are in red, while the lowest across all models are underlined. The results indicate that extrapolating all ranks leads to the worst performance.
@@
6 $1.25$ $1.79$ $2.09$ $2.33$ $1.51$ $1.83$
$1.5$ $1.63$ $1.83$ $1.97$ $1.40$ $1.66$
Full Rank (ExPO) $1$ $1.60$ $1.90$ $2.10$ $1.40$ $1.69$
$1.25$ $1.53$ $1.80$ $1.93$ $1.35$ $1.62$
@@
Figure 8: Harmful request completions for LoX (blue) and ExPO (red).
To measure the degradation effects of extrapolation and show how LoX mitigates this issue, when compared to ExPO, in Table 8 we present utility metrics for the Llama models aligned with 22.5k and 65.6k examples. We compare their utility metrics before extrapolation (base) and after applying ExPO and LoX, with $\alpha=1.25$. The results show a degradation across all five metrics when comparing both LoX and ExPO to the base model, with LoX mitigating this degradation relative to ExPO. This indicates that using only top-ranks is capable of preserving more the original performance of the model, when compared to full-rank.
@@ Appendix G Safety Landscape Details
In this section, we provide auxiliary details on how we constructed the safety landscape plots in Section 5.2.
@@
where $\text{ASR}(\cdot)$ represents the Attack Success Rate of the model, $\alpha,\beta\in\mathbb{R}$, and $d_{1},d_{2}\in\mathbb{R}^{d}$ are two chosen directions.
@@ Appendix H ASR Evaluation Method and Prompts
We use GPT-4o mini to perform the evaluation on the first 100 examples from AdvBench. Responses are generated using greedy decoding from the evaluated model. Each output is rated on a safety scale from 1 to 5, with higher scores indicating more unsafe responses, as defined by the scoring guidelines. The Attack Success Rate (ASR) is calculated as the proportion of responses that receive a score of 5. For the scoring guidelines, we follow (He et al., 2024), which is a revised version from (Qi et al., 2023). We include Meta’s usage guidelines [^1] https://ai.meta.com/llama/use-policy/ in our prompt, following the evaluation protocol of Qi et al. (2023). These guidelines serve to encode policy knowledge into the model, clarifying which uses are acceptable and which are considered harmful. This helps the model better assess the safety of generated completions. The prompt used is presented in Fig. 9.

### v3

@@ Appendix C Effecive Rank Experiments
[…]
@@
[…]
@@
As observed in Fig.˜7, the safety can be recovered by preserving just a few ranks, in all cases. For the Llama model aligned with 65.6k examples, the effective rank is 6, while the other models the effective rank is 3.
@@ Appendix D Extended Robustness Metric Results
In Table˜3, we present individual value results for Fig.˜2 . It is possible to osberve that (a) robustness increases as the data size increases and (b) $R_{\text{ft}}<R_{\text{align}}$, indicating that the proportion of safety knowledge decreases after fine-tuning.
@@ Appendix E Extended Identity Shifting Attack Results
Table˜4, Table˜5, and Table˜6 present individual run results for the experiments in Table˜1, Table˜2, and Fig.˜5, respectively.
Table 4: ASR of individual Identity Shifting attack runs in correspondence to Table˜1.
@@
Table 5: ASR of individual Identity Shifting attack runs in correspondence to Table˜2.
@@
Table 6: ASR of individual Identity Shifting attack runs in correspondence to Fig.˜5.
@@ Appendix F Broken Output Examples
To further showcase the model degradation effects of ExPO when compared to LoX, and quantitatively justify the observation in Section˜5.2, we provide evaluation of the highest $\alpha$ models plotted on Fig.˜5, and extrapolating further (which we considered broken), following (Lin et al., 2023). Table˜7 shows that extrapolating full rank to $\alpha=1.25$ causes the largest degradation according to all metrics. In Fig.˜8, we also present examples of broken and normal completions generated by these models.
Table 7: Utility metrics for $k=6$, full rank, and varying $\alpha$. For both values of $k$, we show the highest $\alpha$ from Fig.˜5 and extrapolating further (which we consider broken). The lowest values for each $k$ are in red, while the lowest across all models are underlined. The results indicate that extrapolating all ranks leads to the worst performance.
@@
6 1.251.251.25 1.791.791.79 2.092.092.09 2.332.332.33 1.511.511.51 1.831.831.83
1.51.51.5 1.631.631.63 1.831.831.83 1.971.971.97 1.401.401.40 1.661.661.66
Full Rank (ExPO) 111 1.601.601.60 1.901.901.90 2.102.102.10 1.401.401.40 1.691.691.69
1.251.251.25 1.531.531.53 1.801.801.80 1.931.931.93 1.351.351.35 1.621.621.62
@@
Figure 8: Harmful request completions generated by LoX (blue) and ExPO (red). ExPO yields broken outputs at $\alpha=1.25$, while LoX remains coherent at this level and only begins to break at $\alpha=1.5$. These examples support the claim that LoX enables extrapolation to higher values of $\alpha$ than ExPO.
To measure the degradation effects of extrapolation and show how LoX mitigates this issue, when compared to ExPO, in Table˜8 we present utility metrics for the Llama models aligned with 22.5k and 65.6k examples. We compare their utility metrics before extrapolation (base) and after applying ExPO and LoX, with $\alpha=1.25$. The results show a degradation across all five metrics when comparing both LoX and ExPO to the base model, with LoX mitigating this degradation relative to ExPO. This indicates that using only top-ranks is capable of preserving more the original performance of the model, when compared to full-rank.
@@ Appendix G Safety Landscape Details
In this section, we provide auxiliary details on how we constructed the safety landscape plots in Section˜5.2.
@@
[…]
@@ Appendix H ASR Evaluation Method and Prompts
We use GPT-4o mini to perform the evaluation on the first 100 examples from AdvBench. Responses are generated using greedy decoding from the evaluated model. Each output is rated on a safety scale from 1 to 5, with higher scores indicating more unsafe responses, as defined by the scoring guidelines. The Attack Success Rate (ASR) is calculated as the proportion of responses that receive a score of 5. For the scoring guidelines, we follow (He et al., 2024), which is a revised version from (Qi et al., 2023). We include Meta’s usage guidelines [^1] https://ai.meta.com/llama/use-policy/ in our prompt, following the evaluation protocol of Qi et al. (2023). These guidelines serve to encode policy knowledge into the model, clarifying which uses are acceptable and which are considered harmful. This helps the model better assess the safety of generated completions. The prompt used is presented in Fig.˜9.